OpenStudy (anonymous):

Which is the ewuation for the line perpendicular to y= -5/3x + 11 1/3 and conataining p(-2,3)? Hellpp y- 2 = -3/5(x-3) y=-5/3x+4 1/5 y = -3/5x + 4 1/5 y= 3/5x + 4 1/5

OpenStudy (anonymous):

From our last problem, you take the negative reciprocal of the slope of the line in the problem statement. Can you get that?

OpenStudy (anonymous):

i have no idea what u just said

OpenStudy (anonymous):

Then you use the point-slope form for the equation of a line.\[y - y _{1} = m(x - x _{1})\] where "m" is the new slope.

OpenStudy (anonymous):

Ok, we'll start with the line in the problem. It is in the slope-intercept form which is y = mx + b where "m" is the slope. Just match up the positions. and you will get the slope easily.

OpenStudy (anonymous):

would it be -5/3 honestly i have no idea

OpenStudy (anonymous):

-5/3 is the slope of the equation in the problem. You did a good job of identifying it. So, the slope we want for our new line is 3/5 which is the negative reciprocal, which is switching the numerator and denominator and making it negative.